OpenStudy (sonyalee77):

What is the equation of the line that passes through the points [1/2,-1] and (1,1)?

OpenStudy (noseboy908):

are you familiar with the equation for slope?

OpenStudy (sonyalee77):

y=mx+b?

OpenStudy (noseboy908):

no, that's slope-intercept form.

OpenStudy (noseboy908):

so, do you know the formula for slope?

OpenStudy (sonyalee77):

no

OpenStudy (noseboy908):

ok, it's (y2-y1)/(x2-x2)

OpenStudy (sonyalee77):

oh now i remember thanx!!!

OpenStudy (noseboy908):

so you know what to do now

OpenStudy (noseboy908):

??

OpenStudy (sonyalee77):

more or less

OpenStudy (noseboy908):

what is still confusing? (even to a small degree)

OpenStudy (sonyalee77):

my answer was y=1/2 but it was not one of my answer options

OpenStudy (noseboy908):

Please elaborate. Y=1/2 is a horizontal line.

OpenStudy (sonyalee77):

A. y = 1 B. y=1/4x+3/4 C. y = 4x + 3 D. y = 4x – 3

OpenStudy (noseboy908):

Firstly, tell me how you found the slope.

OpenStudy (sonyalee77):

1-(-1)/1-1/2

OpenStudy (noseboy908):

it's upside down, you did x2-x1 on top. it should be on the bottom.

OpenStudy (noseboy908):

\[(y2-y1)/(x2-x1)\]

OpenStudy (sonyalee77):

dont get it

OpenStudy (sonyalee77):

@superdavesuper

OpenStudy (noseboy908):

ok, take your given two points: (1/2,-1) and (1,1). Think about what an ordered pair is: an x, paired with a y, and restated as "(x,y). So, take the second given point, and it becomes (x2,y2). Whereas the first pair stays as (x1,y1)

OpenStudy (noseboy908):

so put the two y's, and the two x's into the formula I gave you.

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ \frac{1}{2}}}\quad ,&{\color{blue}{ -1}})\quad &({\color{red}{ 1}}\quad ,&{\color{blue}{ 1}}) \end{array} \\\quad \\ slope = {\color{green}{ }}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ y-y_1={\color{green} m}(x-x_1)\quad \textit{plug in your values and solve for "y"}\)

OpenStudy (jdoe0001):

keep in mind that for the point-slope form, that is \(\bf y-y_1={\color{green} m}(x-x_1)\) you can use either point, \(\bf (x_1,y_1) \ or \ (x_2,y_2)\)

OpenStudy (sonyalee77):

is it a y=1? tht was my answer

OpenStudy (jdoe0001):

dunno.. what did you get for the slope? \(\bf slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}}\)

OpenStudy (sonyalee77):

1/2

OpenStudy (sonyalee77):

y intercept was 2

OpenStudy (sonyalee77):

D?

OpenStudy (noseboy908):

Show me your process for finding the slope.

OpenStudy (jdoe0001):

got a bit .lagged... one sec

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ \frac{1}{2}}}\quad ,&{\color{blue}{ -1}})\quad &({\color{red}{ 1}}\quad ,&{\color{blue}{ 1}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}}\implies \cfrac{{\color{blue}{ 1}}-{\color{blue}{ (-1)}}}{{\color{red}{ 1}}-{\color{red}{ \frac{1}{2}}}}\)

OpenStudy (jdoe0001):

\(\bf \cfrac{{\color{blue}{ 1}}-{\color{blue}{ (-1)}}}{{\color{red}{ 1}}-{\color{red}{ \frac{1}{2}}}}\implies \cfrac{2}{\frac{1}{2}}\\ \quad \\ \textit{recall that }\cfrac{\quad \frac{a}{b}\quad }{\frac{c}{d}}\implies \cfrac{a}{b}\cdot \cfrac{d}{c}\qquad thus\\ \quad \\ \cfrac{2}{\frac{1}{2}}\implies \cfrac{\frac{2}{1}}{\frac{1}{2}}\implies \cfrac{2}{1}\cdot \cfrac{2}{1}\)